Paul's age is two years more than twice Mark's age. The sum of their ages is 32 years. How old are Paul and Mark?
Answer by lenny460(1073)

Let Mark's age be n
Let Paul's age be n * 2 + 2 = 2n + 2
Combine their ages:
n + 2n + 2 = 32
Combine like terms:
n + 2n = 3n
3n + 2 = 32
Subtract 2 from each side of the equation
3n + 2 - 2 = 32 - 2
Combine like terms:
2 - 2 = 0
32 - 2 = 30
3n = 30
Divide each side by 3
3n/3 = 30/3
n = 10
2n + 2 = 2 * 10 + 2
= 20 + 2 = 22
Therefore:
Mark is 10 years old and
Paul is 22 years old.